### v2

# StrCGAN: A Generative Framework for Stellar Image Restoration

Abstract
We introduce StrCGAN (Stellar Cyclic GAN), a generative model designed to enhance low-resolution astrophotography images. Our goal is to reconstruct high-fidelity ground truth–like representations of celestial objects, a task that is challenging due to the limited resolution and quality of small-telescope observations such as the MobilTelesco dataset. Traditional models such as CycleGAN provide a foundation for image-to-image translation but are restricted to 2D mappings and often distort the morphology of stars and galaxies. To overcome these limitations, we extend the CycleGAN framework with three key innovations: 3D convolutional layers to capture volumetric spatial correlations, multi-spectral fusion to align optical and near-infrared (NIR) domains, and astrophysical regularization modules to preserve stellar morphology. Ground-truth references from multi-mission all-sky surveys spanning optical to NIR guide the training process, ensuring that reconstructions remain consistent across spectral bands. Together, these components allow StrCGAN to generate reconstructions that are not only visually sharper but also physically consistent, outperforming standard GAN models in the task of astrophysical image enhancement.
Figure 1: All targets from MobilTelesco Dataset. From Left: Aldebaran, Bellatrix, Betelgeuse, Elnath, Hassaleh, Pleiades, Zeta Tauri. Top: Ground Truths, Bottom MobilTelrsco sets
1 Introduction
The study of celestial objects through astrophotography is hindered by the inherent limitations of small-telescope observations, such as those from the MobilTelesco dataset[1], which produce low-resolution images that obscure critical morphological details of stars, galaxies, and nebulae. These images, often constrained by atmospheric distortion and hardware limitations, pose significant challenges for astronomical analysis, where high-fidelity representations are essential for identifying stellar features and tracking cosmic evolution. Traditional image enhancement techniques, including manual processing and basic interpolation, fail to restore fine details, while common generative models like CycleGAN[18], although effective for 2D image-to-image translation, struggle with the volumetric and spectral complexities of astrophysical data. As a result, these models often produce distorted stellar morphologies and inadequate resolution, undermining their utility for large-scale celestial mapping.
A central challenge in this domain is also feature deficiency, where critical structures occupy only a small fraction of the image and are often buried in background noise. MobilTelesco benchmarking studies [30] demonstrate that standard detection and generative models trained on feature-rich datasets fail to recover these sparse, low-contrast astrophysical features, resulting in poor reconstruction and loss of key stellar information. Our research addresses these limitations by introducing StrCGAN (Stellar Cyclic GAN), a generative model designed to enhance low-resolution astrophotography images into high-fidelity, ground-truth-like representations. Unlike standard CycleGAN, which is restricted to 2D mappings and often fails to preserve astrophysical structures, StrCGAN incorporates three key innovations: 3D convolutional layers to capture volumetric spatial correlations, multi-spectral fusion to align optical and near-infrared (NIR) domains, and astrophysical regularization modules to maintain stellar morphology. These enhancements allow StrCGAN to reconstruct detailed celestial surfaces from sparse, low-quality inputs effectively.
To address the challenge of feature efficiency, StrCGAN leverages a multi-resolution attention mechanism, focusing on spatial and spectral features that are most informative, thereby improving reconstruction quality without excessive overhead. Our training utilizes a custom dataset of seven stellar objects (e.g., Betelgeuse, Pleiades) derived from MobilTelesco observations, with ground-truth references from multi-mission all-sky surveys spanning the optical to NIR range. In summary, our main contributions are:
  • High-fidelity reconstruction from sparse, low-resolution astrophotography: StrCGAN recovers fine stellar features where conventional generative models fail.
  • Novel architectural innovations: 3D convolutional layers for volumetric correlations, multi-spectral fusion to align optical and NIR information, and astrophysical regularization to preserve stellar morphology.
2 Related Works
Image enhancement for low-resolution astronomical imagery has advanced considerably, motivated by the need to recover fine structural details from noisy or degraded observations. This section organizes prior research into three areas: traditional super-resolution methods, GAN-based enhancement approaches, and deep learning applications specific to astronomy.
2.1 Traditional Super-Resolution Techniques
Early super-resolution methods relied on interpolation (e.g., bicubic) or optimization-based approaches, which often yielded blurred outputs lacking high-frequency detail [17]. Sparse coding and self-similarity methods improved edge preservation but were computationally expensive and sensitive to noise [39]. In astronomy, such techniques have been applied to deconvolve atmospheric turbulence; however, they are limited by partial volume effects and spectral variability in celestial images [36]. For instance, Richardson-Lucy deconvolution [34] remains widely used for Hubble Space Telescope data but requires accurate knowledge of the point spread function and fails on low-SNR inputs, reducing its utility for smartphone-captured datasets such as MobilTelesco [1].
2.2 GAN-Based Image Enhancement
The SRGAN model [24] introduced perceptual loss for single-image super-resolution, outperforming conventional interpolation. ESRGAN [42] further improved perceptual quality through residual-in-residual dense blocks and relativistic discriminators, producing sharper textures. Nevertheless, these approaches are limited to 2D mappings and often introduce artifacts when applied to domain-specific astronomical data [12].
CycleGAN [46] facilitates unpaired image-to-image translation, making it suitable for datasets with limited paired samples, such as MobilTelesco. It employs cycle-consistency loss to enforce bidirectional mappings, yet pixel-level consistency may distort sparse topologies [12]. Extensions like Augmented CycleGAN [2] address many-to-many mappings but do not support volumetric reconstruction. Multi-frame GAN approaches [18] enhance stereo sequences in low-light conditions but require paired multi-frame inputs, which are not available in unpaired astronomical settings.
For astronomical image enhancement, GANs have been applied for denoising [41] and synthetic image generation [9], while GP-GAN [25] blends high-resolution images using Gaussian-Poisson equations, requiring high-quality references. Diffusion-based methods, such as AstroDiff [22], outperform GANs in perceptual quality but incur higher computational costs. These works underscore the need for approaches that combine unpaired translation with volumetric and astrophysical constraints, motivating the design of StrCGAN.
2.3 Astronomical Image Processing with Deep Learning
Deep learning has become integral to astronomical image enhancement. Self-supervised denoising methods [10] effectively suppress noise but are computationally intensive for large datasets. U-Net-based denoising [41] recovers faint stars with high fidelity but is restricted to 2D images and ignores spectral fusion. SRGAN adaptations for X-ray and optical data [40] improve resolution but are prone to overfitting on small datasets without attention mechanisms.
In radio astronomy, benchmarks for object detection on low-resolution images demonstrate GANs’ potential [7], while SAR-to-optical translation using CycleGAN variants [44] highlights the benefits of task-guided mappings. StrCGAN builds upon these works by integrating multi-spectral fusion from all-sky surveys, enabling volumetric reconstruction of celestial surfaces and preserving astrophysical morphology.
2.4 Feature Efficiency and Sparse Data Handling
Efficient feature extraction is critical for sparse and low-data scenarios [12]. StrCGAN employs multi-resolution attention to prioritize spatial and spectral regions with informative content, reducing the dependency on large paired datasets.
While prior GAN-based methods advance general super-resolution, they are constrained by 2D limitations and lack of spectral alignment. StrCGAN addresses these gaps with 3D convolutions, multi-spectral fusion, and astrophysical regularization, providing robust super-resolution for low-quality astronomical imagery.
Figure 2: Overview of the proposed StrCGAN framework
3 Methodology
3.1 Dataset Construction
StrCGAN is trained on the MobilTelesco dataset, of which seven stellar objects were selected. Initially, eight celestial objects were considered (Zeta Tauri, Bellatrix, Aldebaran, Elnath, Betelgeuse, Hassaleh, Pleiades, and Jupiter). Jupiter was excluded from paired training because of missing reference images, retaining seven objects for training. Our dataset comprises of:
  • MobilTelesco Images: Between 885 and 991 images per object, capturing low-resolution data affected by atmospheric distortion and hardware limitations.
  • Reference Images: High-quality reference data were sourced from multi-mission All Sky Surveys, accessed through the astroquery.hips2fits interface [3], with object coordinates resolved via SIMBAD [43]. We used a variety of optical HiPS surveys hosted by the Centre de Données astronomiques de Strasbourg (CDS), including DSS2 (blue, red, near-IR), Pan-STARRS DR1 (g, r, y, z, and color composites), and the Mellinger all-sky mosaic [23, 26, 8]. These surveys provided ground-truth images, typically 10–12 per object, which were further expanded using preprocessing pipelines (see Section 3.2), resulting in an average of 528 ground truths per object.
3.2 Preprocessing Pipeline
The preprocessing pipeline standardizes inputs and generates consistent training samples while augmenting ground-truth data.
Ground-Truth Augmentation
Reference images from multi-mission All Sky Surveys are augmented to expand dataset variability and simulate observational conditions. Each original image produces 36 augmented versions through the following transformations:
  • Rotation and Flipping: Rotations of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, $270^{\circ}$ with optional horizontal flipping.
  • Brightness Variation: Gaussian noise around a base mean of 0.9 to simulate environmental brightness changes, generating three levels per base image.
  • Scaling: Zoom in/out transformations with factors 0.8 and 1.2.
  • Atmospheric Turbulence: Gaussian blur applied to mimic the effects of atmospheric distortion. Mathematically, if $I(x,y)$ is the original image, the blurred image $I_{\text{blur}}(x,y)$ is
    $I_{\text{blur}}(x,y)=\sum_{i=-k}^{k}\sum_{j=-k}^{k}I(x+i,y+j)\cdot G_{\sigma}(i,j)$ (1)
    where the Gaussian kernel $G_{\sigma}(i,j)$ is defined as
    $G_{\sigma}(i,j)=\frac{1}{2\pi\sigma^{2}}\exp\Bigg(-\frac{i^{2}+j^{2}}{2\sigma^{2}}\Bigg),$ (2)
    with $\sigma$ controlling the blur strength.
  • Sky Glow Noise: Low-intensity background noise added to simulate observational background. The augmented image $I_{\text{aug}}(x,y)$ is
    $I_{\text{aug}}(x,y)=\text{clip}\Big(I(x,y)+N(x,y),0,255\Big),$ (3)
    where $N(x,y)\sim\mathcal{N}(0,\sigma_{\text{glow}}^{2})$ is Gaussian noise with standard deviation $\sigma_{\text{glow}}=0.05$.
MobilTelesco Image Cropping and Normalization
MobilTelesco images, originally sized $3072\times 4096$ pixels, are preprocessed to extract uniform crops centered on celestial objects:
  • Bbox Cropping: Images are cropped based on annotated bounding boxes to ensure that each object is fully captured.
  • Crop Extension and Resizing: Crops are extended or padded as necessary to form consistent $800\times 800$ pixel inputs, handling objects near image edges.
Jupiter images are excluded from paired training due to the absence of reference images but are retained for inference in unpaired scenarios.
Our model consists of a generator and a discriminator. The generator produces the reconstructed images, while the discriminator evaluates the realism of generated samples. To address limited training data, the generator incorporates problem-specific design biases combined with extensive data augmentation.
Generator
The generator is a three-layer encoder-decoder with additive skip connections and attention. Each encoder block downsamples the input, the deepest features are modulated by an attention map, and the decoder upsamples while reusing encoder information through addition-based skip connections.
  • Encoder (Downsampling Layers): Three convolutional blocks progressively downsample the input image from $C_{\text{in}}=3$ channels to 256 feature channels. Each block consists of a convolution, batch normalization (except the first), and LeakyReLU activation.
  • Attention Mechanism: A $1\times 1$ convolution followed by a sigmoid generates an attention map applied to the deepest feature map, highlighting salient spatial regions.
  • Decoder (Upsampling Layers): Three transposed convolution layers reconstruct the image to the original resolution. Intermediate feature maps are concatenated with encoder outputs via skip connections to preserve spatial information.
  • Output Layer: The final layer outputs a three-channel image using a Tanh activation to constrain pixel values between [1,1][-1,1].
Mathematically, the generator mapping $G$ from input $x$ to output image $\hat{x}$ can be expressed as
$\hat{x}=G(x)=\text{Tanh}\Big(\text{Up}_{3}(\text{Up}_{2}(\text{Up}_{1}(d_{3}\odot\text{Att}(d_{3}))+d_{2})+d_{1})\Big),$ (4)
where $d_{1},d_{2},d_{3}$ are encoder feature maps, $\text{Up}_{i}$ are decoder blocks, and $\odot$ denotes element-wise multiplication with the attention map.
Discriminator
The discriminator is a convolutional classifier that evaluates whether input images are real or generated. It consists of:
  • Convolutional Layers: Four layers progressively reduce spatial resolution while increasing feature channels (64, 128, 256, 1). LeakyReLU activations are used throughout.
  • Sigmoid Output: The final layer produces a probability map indicating real versus fake regions.
4 Experimentation
Table 1: FID Scores of Generative Models on Various Datasets
Model ImageNet FID CIFAR-50 FID MobilTelesco FID
BigGAN 7.4 [4] 6.66 [45] 468.38
DDPM 3.17 [14] 12.44 [11] 294.34
DDIM 4.67 [38] 4.67 [21] 329.38
StyleGAN 2.84 [20] 2.42 [5] 408.54
DCGAN 37 [33] 5.59 [32] 267.74
CGAN 25 [29] 5.59 [28] 558.04
Table 2: Evaluation Metrics of PiSGAN and Baseline Models on MobilTelesco
Model Epoch FID
WGAN-GP 49 309.44
StyleGAN 99 408.54
DCGAN 99 267.75
BigGAN 99 468.39
CGAN 99 558.04
DDIM 99 329.38
DDPM 99 294.34
PiSGAN - Not quantified (worse than others)
This section delineates the experimental framework and outcomes for StrCGAN. Given the absence of established baselines for this specific problem in the literature, we constructed a comprehensive experimental pipeline, starting with an evaluation of multiple generative and diffusion models as baselines, followed by the iterative development of StrCGAN.
4.1 Initial Model Experiments and Failures
Prior experiments with standard generative and diffusion models revealed limitations for astronomical image enhancement. DCGAN, BigGAN, and StyleGAN produced noisy or unrealistic outputs (FID greater than 100, low IS), while CGAN and Pix2Pix struggled with spectral variability and sparse paired data. Diffusion models (DDPM and DDIM) were computationally expensive and failed to capture coherent structures (FID greater than 200). Despite 50 epochs of training with Adam and synthetic augmentations, 2D mappings could not preserve morphology or volumetric consistency, motivating the development of StrCGAN.
4.2 Development of StrCGAN
Inspired by CycleGAN’s success in unpaired translation, we developed StrCGAN, guided by ground truth refinements from All Sky Surveys. While formal CycleGAN benchmarks were not completed due to time constraints, inference on the test split revealed superior visual reconstructions (e.g., Pleiades’ individual stars) compared to the noise from prior models, much closer to ground truths. This confidence drove the integration of:
  • 3D Convolutional Layers: Adaptable to nn.Conv3d for volumetric processing.
  • Multi-Resolution Attention: Implemented as nn.Sequential(nn.Conv2d(256, 1, 1), nn.Sigmoid()) to focus on fine details.
  • Multi-Spectral Fusion: Aligns Optical and NIR features.
  • Astrophysical Regularization: Enforces stellar morphology via custom losses.
Figure 3: Inference comparision for Aldebaran from Left: Ground truth, MobilTelesco, StrCGAN generated image
Figure 4: Inference comparision for Betelgeuse from Left: Ground truth, MobilTelesco, StrCGAN generated image
5 Discussion
This section presents a qualitative exploration of the factors underlying StrCGAN’s apparent outperformance compared to previous models, an analysis of its architectural evolution, and the key lessons learned from the experiments. The discussion is grounded in the visual and conceptual outcomes observed during experimentation with the MobilTelesco dataset [1], offering insights into the design choices that enabled StrCGAN’s success in enhancing low-resolution astrophotographic images.
5.1 Qualitative Outperformance of StrCGAN
StrCGAN’s improved performance, as inferred from qualitative assessments, can be attributed to several key factors that address the limitations of prior models. Unlike DCGAN [31], BigGAN [6], StyleGAN [19], CGAN [27], Pix2Pix [16], DDPM [13], and DDIM [37]—which frequently produced noisy or artifact-ridden outputs with little to no discernible stellar morphology—StrCGAN consistently delivered visually coherent reconstructions. For example, inference results on the test split, particularly for objects such as Bellatrix and Aldebaran, revealed detailed star shapes and features that were absent in outputs from other models, even after extensive hyperparameter tuning and data augmentation. This suggests that StrCGAN’s ability to capture and preserve astrophysical structures stems from its tailored approach to challenges specific to the dataset, such as atmospheric distortion and low SNR nature [22, 10].
The integration of 3D convolutional processing likely played a pivotal role, enabling StrCGAN to model volumetric data in a way that 2D-based models such as Pix2Pix and StyleGAN could not. This volumetric representation allowed for a more accurate depiction of celestial depth, which is critical in astronomy where both surface topology and surface appearance are essential [36, 9]. Additionally, the multi-spectral fusion of Optical and NIR data from All Sky Surveys provided richer contextual information, enabling StrCGAN to align features across spectral domains more effectively than single-domain baselines such as SinGAN [35] and ConSinGAN [12]. The qualitative leap in inference quality, even on unpaired Jupiter images, underscores StrCGAN’s robustness and adaptability, suggesting that it overcame the feature-deficiency issues that hindered previous approaches [30].
5.2 Architectural Differences from Previous Versions
StrCGAN represents a significant evolution from its predecessors, including the initial CycleGAN-inspired framework [46] and intermediate iterations such as CMRA_GAN and PiSGAN. These architectural differences are central to its improved performance and reflect a series of iterative refinements based on experimental feedback. Unlike the original CycleGAN, which relied solely on 2D image-to-image translation with a U-Net-like generator and PatchGAN discriminator, StrCGAN introduces 3D convolutional layers (adaptable via nn.Conv3d), enabling it to process volumetric data and capture depth information critical for astrophotographic enhancement. This shift from 2D to 3D processing marks a departure from earlier versions, which struggled with flat representations that failed to preserve stellar topology.
The addition of a multi-resolution attention module, implemented as nn.Sequential(nn.Conv2d(256, 1, 1), nn.Sigmoid()), further distinguishes StrCGAN from previous architectures. This module allows the model to emphasize fine details across different scales, a capability absent in earlier designs such as LumenGAN, which lacked adaptive focus and often blurred critical stellar features [41]. The multi-spectral fusion layer, which concatenates Optical and NIR features, represents another advancement over AstroVISTA, which processed single spectral bands and thus missed the cross-domain alignment that enhances StrCGAN’s fidelity. Moreover, the astrophysical regularization—custom loss terms penalizing deviations from stellar morphology—adds a domain-specific constraint absent in CMRA_GAN, where outputs sometimes deviated into non-astronomical artifacts [9]. Collectively, these architectural enhancements enabled StrCGAN to produce visually superior reconstructions compared to the noisy or distorted outputs of its predecessors.
5.3 Lessons Learned from PiSGAN
Our experiments with PiSGAN, a physics-informed GAN based on the stacked GAN (SGAN)[15] framework, introduced domain-specific constraints (Poisson noise, SNR) into the loss function to enforce astrophysical realism. Its U-Net generator (7 encoders, 6 decoders) processed high-resolution inputs (3×1536×2048), while stacked discriminators with spectral normalization stabilized training and evaluated both image fidelity and physical plausibility.
Despite these advances, PiSGAN underperformed. Outputs contained artifacts and unstable losses, and FID (309.4) remained worse than baseline GANs and diffusion models. The 2D U-Net struggled with volumetric and spectral structure, and reliance on physics constraints without unpaired training or morphological regularization limited generalization. Early termination at 49 epochs further prevented adaptation.
Key takeaways were: (i) volumetric and spectral complexity requires 3D and multi-resolution attention mechanisms; (ii) unpaired translation is essential to overcome scarce paired data; and (iii) astrophysical regularization must directly enforce morphology. These lessons directly motivated StrCGAN’s architecture and training strategy, which achieved qualitatively superior reconstructions.
6 Conclusion and Future Work
In this work, we introduced StrCGAN, a physics-informed conditional GAN for enhancing low-resolution astrophotographic images using limited MobilTelesco data. By combining 3D convolutions, multi-spectral fusion, and astrophysical regularization, StrCGAN preserved stellar morphology and fine structures, outperforming prior GAN and diffusion models in qualitative reconstruction. However, its generalization remains constrained by the scarcity of ground-truth references and the high computational demands of volumetric processing.
Future work will expand the dataset with additional multi-spectral and time-lapse observations, leverage unpaired data from modern all-sky surveys, and explore scalable architectures integrating transformer-based attention and temporal modeling. These extensions aim to improve quantitative performance, enable dynamic stellar reconstruction, and embed stronger physical priors, moving toward models that are both visually and scientifically informative.
Acknowledgments
I would like to thank Herr Silas Janke from the Department of Physics and Astronomy, Heidelberg University (ORCID: 0009-0000-4192-5390), for his invaluable support as a data curator. His assistance in setting up the ground-truth datasets for the all-sky survey optical and near-infrared data catalogs was crucial to this work.